OpenStudy (anonymous):

How do you simplify an expression like the expression 10h + 6 - 5h + 3?

Vocaloid (vocaloid):

put the h's together 10h - 5h = ?

OpenStudy (koalagurl1):

5h+9

OpenStudy (jango_in_dtown):

10h + 6 - 5h + 3=(10-5)h+(6+3)=5h+9

OpenStudy (jango_in_dtown):

a.x+b.x=(a+b)x

OpenStudy (jango_in_dtown):

distributive property

OpenStudy (richgall____):

5h+9

OpenStudy (koalagurl1):

10h-5h=5h

OpenStudy (koalagurl1):

6+3=9

OpenStudy (koalagurl1):

5h+9

OpenStudy (xkshx):

Combine like terms: 10h - 5h = ? and 6+3 = ?

OpenStudy (richgall____):

you have to use to sign in front of the 5h to determine whether to add or subtract then you use the sign in front of them other number to finish the expression

OpenStudy (anonymous):

My question says: Simplify the expression. It has to be a essay and has to be at least 2 sentences.

OpenStudy (anonymous):

Can someone answer my question and give me 2 sentence I promise 100% that I will not use it for my test I will just use it to get an idea of what to write.

OpenStudy (anonymous):

Ok?

OpenStudy (anonymous):

@xkshx

OpenStudy (anonymous):

@jango_IN_DTOWN @KoalaGurl1

OpenStudy (xkshx):

To simplify the following expression you would have to combine like terms. To make this easier you could put the equation descending order: 10h - 6h + 6 + 3. The like terms in the problem are h and coefficients. 10h - 5h equals 5h while 6 + 3 is 9. The simplified equation is 5h +9.

OpenStudy (anonymous):

@richgall____

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

someone answer?

OpenStudy (anonymous):

@xkshx you mean 5h not 6h.

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@N2OPowered

OpenStudy (xkshx):

Yes sorry that was a typo:)

OpenStudy (anonymous):

good.

OpenStudy (anonymous):

well it's not good there was a typo its good that you don't have a problem that caused you to do that.

OpenStudy (anonymous):

xkshx can you right me one in the form of "I got the answer by... The answer is 5h + 9.

OpenStudy (xkshx):

I gave you who to do it, now you have to put it in your own words:)

OpenStudy (anonymous):

No that form I didn't understand @xkshx .

OpenStudy (anonymous):

Can u do it in the form I showed you. That's not how I will do it.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

@xkshx ?????

OpenStudy (anonymous):

Heelo anyone there come on @xkshx I know you are there.

OpenStudy (anonymous):

I can see you in the viewer list.

OpenStudy (anonymous):

Please :(

OpenStudy (anonymous):